# Point-MoE: Towards Cross-Domain Generalization in 3D Semantic Segmentation via Mixture-of-Experts

### v1

@@ B.3 Additional Main Results
Results without precise evaluator. The results reported in Tab. 1 and Tab. 2 are obtained using the precise evaluator provided by Pointcept [20]. The precise evaluator first splits each scene into multiple overlapping chunks, then performs model inference on each chunk individually. After inference, it aggregates predictions across chunks through a voting mechanism to refine the final output. We typically observe a gain of 2–4% in accuracy on each dataset using this method. While effective, the precise evaluator introduces additional runtime overhead during evaluation. Moreover, we note that this form of ensembling may hurt overall performance when the model’s predictions are unreliable in early chunks, as incorrect outputs are also accumulated in the final vote. To support convenient and reproducible benchmarking for the community, we include results with the precise evaluator in the main text and provide additional baseline results without it elsewhere in the paper.
From both Tab. 1 and Tab. 2, we observe that Point-MoE outperforms the baselines in most domains in both the indoor-only and indoor-outdoor joint training settings. These results are consistent with those obtained using the precise evaluator reported in the main text. We do not observe any contradictory trends, as the precise evaluator typically improves all methods by a similar margin of 2–4%. As such, our analysis and conclusions align closely with those presented in the main text. How Strong is Domain-Aware Point-MoE? In the main text, we compare domain-agnostic and domain-aware settings and observe that domain-aware Point-MoE does not consistently outperform its domain-agnostic counterpart. Here we want to offer discussions and analysis to explain the reason. One contributing factor is the reliance on the domain predictor as we detailed in sec. B.2. Although the domain predictor achieves high classification accuracy, even minor errors can lead to noticeable performance degradation—either by misrouting inputs to suboptimal experts in Point-MoE or by selecting incorrect normalization statistics in PPT. To better isolate this effect, we include results in Tab. 12 and Tab. 13, where we evaluate domain-aware Point-MoE using ground-truth domain labels.
In general, Point-MoE continues to outperform standard baselines in both settings. Moreover, we find that domain-aware Point-MoE with access to ground-truth labels achieves better performance on seen domains, while domain-agnostic Point-MoE trained with domain-randomized data generalizes better to unseen domains. Interestingly, we also observe that Point-MoE trained with generic (i.e., domain-agnostic) supervision performs well even on seen domains when exposed to more diverse training inputs. This suggests that domain-randomization during training may benefit performance across both seen and unseen domains in challenging settings. Finally, we note that domain-aware Point-MoE can yield stronger results when accurate domain labels are available at inference time.
@@ Appendix C Additional Visualization and Discussions on Point-MoE
C.1 t-SNE
Figure 6: t-SNE Visualization. The first three columns illustrate feature similarities and differences across PTv3, PPT, and Point-MoE for both encoder and decoder layers. The rightmost column highlights how Point-MoE generalizes to unseen domains, with Matterport3D features aligning closely with the most relevant seen domains—demonstrating robust cross-domain generalization.
Fig. 6 presents t-SNE visualizations [69] of encoder and decoder features from PTv3, PPT [80], and Point-MoE. PTv3 exhibits no clear separation in either encoder or decoder spaces, indicating its inability to structure domain-specific representations. In contrast, PPT shows strongly separated clusters for each domain at both encoder and decoder levels, suggesting it overfits to domain-specific patterns. While this benefits seen domains, it may hinder generalization to unseen settings.
Point-MoE, however, reveals a more interesting behavior. Its encoder features remain entangled across domains—indicating shared representation learning—while its decoder learns to disentangle domain-specific structures. This division of labor suggests that Point-MoE effectively performs domain inference within the model, without requiring explicit domain labels. Most notably, in the unseen Matterport3D domain, Point-MoE maps samples to nearby clusters of the most relevant seen domains (e.g., ScanNet), enabling strong generalization. This implicit alignment underscores the model’s ability to organize and interpret new environments by leveraging expertise learned from diverse sources.
C.2 Semantic Concept Distribution over Experts
@@
C.4 More Expert Choice Visualization

### v2

@@ B.3 Additional Main Results
Results without precise evaluator.
The results reported in Tab. 1 are obtained using the precise evaluator provided by Pointcept [18]. The precise evaluator first splits each scene into multiple overlapping chunks, then performs model inference on each chunk individually. After inference, it aggregates predictions across chunks through a voting mechanism to refine the final output. We typically observe a gain of 2–4% in accuracy on each dataset using this method. While effective, the precise evaluator introduces additional runtime overhead during evaluation. Moreover, we note that this form of ensembling may hurt overall performance when the model’s predictions are unreliable in early chunks, as incorrect outputs are also accumulated in the final vote. To support convenient and reproducible benchmarking for the community, we include results with the precise evaluator in the main text and provide additional baseline results without it elsewhere in the paper.
From both Tab. 12 and Tab. 13, we observe that Point-MoE outperforms the baselines in most domains in both the indoor-only and indoor-outdoor joint training settings. These results are consistent with those obtained using the precise evaluator reported in the main text. We do not observe any contradictory trends, as the precise evaluator typically improves all methods by a similar margin of 2–4%. As such, our analysis and conclusions align closely with those presented in the main text. To ensure a strong baseline, we add mixed-dataset training and tune hyper-parameters, reporting the baseline with its best configuration. To be specifically, we apply mix-dataset training for baselines we can directly apply without code modification and we also find layernorm is empirically better for the baselines.
@@ Appendix C Additional Visualization and Discussions on Point-MoE
C.1 Semantic Concept Distribution over Experts
@@
C.3 More Expert Choice Visualization